OpenStudy (anonymous):

Find the value by referring to the graph of the sine function. sin(-7/2 pi)

OpenStudy (anonymous):

\[\sin (-\frac{ 7 }{ 2 }\pi)\]

OpenStudy (idku):

be back....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I don't think so. I think it wants me to use the graph of the function